# Dance to Demise – How Massive Stars May Form Dense Circumstellar Shells Before Explosion

### v2

@@ 6.4 Implications for pre-explosion dust formation around RSGs
RSG stars are often associated with a dusty circumstellar medium (J. J. Walmswell & J. J. Eldridge, 2012; G. C. Clayton et al., 2009; E. Cannon et al., 2021; Z. Niu et al., 2023; G. Folatelli et al., 2025). Yet the modest mass loss rates ( 10-6 M⊙) characterizing RSG winds are found unsuitable to synthesize dust grains from the gas-phase atoms (I. Cherchneff, 2013). Dust formation requires moderately low temperatures while the gas densities are still high to allow series of collisions leading to the formation of gas-phase precursors (A. Sarangi et al., 2018). In the RSG with a dense CSM, the enhanced mass loss can aid the formation of chemical compounds in the wind, depending on the wind velocities. As we have shown, each stellar model evolves in unique routes, which will result in a unique distribution of molecules and dust grains. If the enhanced mass loss is triggered during the He-burning phase or early C-burning phase, the dust formed in that gas will fly far from the star by the time the star goes to SN. This dust may already integrate with the ambient interstellar medium (ISM), and enhance the dust budget in the galaxy. Moreover, given the large distance, they are shielded from the strong ionizing SN flashwave at the point of core-collapse. However, this dust is difficult to detect in stellar spectroscopy in optical or IR, given it will be at a large distance (hence causing no significant extinction to the stellar light curve) and cold temperatures (making it indistinguishable with the ambient ISM).
On the other hand, in the RSGs with enhanced mass loss at the very end of their lives, the dusty CSM significantly impacts the stellar light curve and its post-explosion IR signatures. In addition, the composition of the CSM may also be altered based on the nature of the stellar evolution and mass loss. Even though the CSM is often identical in composition with the H-rich stellar envelope, a small variation in the abundance of the metals will impact the composition and dust masses. The formation of dust precursors in the gas is often controlled by some bottleneck gas-phase reactions (A. Sarangi & I. Cherchneff, 2013). On the contrary, the growth of grains, which decides the grain size distributions, is proportional to time (H. Hirashita, 2012). The population of larger grains are more resilient to destruction processes post SN explosion, or shocks in the ISM (E. R. Micelotta et al., 2018). In that regard, the timing of the enhanced mass loss plays key role in determining the grain sizes.
@@ 6.4.1 Numerical and computational challenges in obtaining pre-SN yields
In order to predict pre-explosion yields to obtain dust masses, our MESA runs need to be evolved through explosive oxygen and silicon burning that occurs on day-hour timescales. As already discussed, accurate energy generation requires the use of large nuclear networks (e.g. mesa_128.net) with implicit hydrodynamics and timesteps causing severe slowdowns (with a single run requiring more than a week on a -core CPU machine), making a similar systematic exploration of our parameter space in and , prohibitively expensive. A hybrid computing strategy – involving zone-by-zone parallelism on a GPU by offloading the nuclear network calculations that form a major bottleneck of MESA with large nuclear networks – seems a promising approach to significantly reduce the large computational cost of evolving our runs through to CC. With an initial grid of such models, the recently proposed Nuclear Neural Network (NNN) framework of A. Grichener et al. (2025) can be employed to further refine the grid that seems essential to study stochastic effects of convection, shell merging etc. on final yields and associated dust masses.
@@ 7 Conclusions
This work presents a comprehensive, end-to-end investigation of how pulsating RSGs can create a dense, possibly dusty, CSM through enhanced mass loss episodes in the final years before CC, which shapes the multi-wavelength appearance of the ensuing SN explosion. From a dense grid of MESA models of varying M⊙ and PDSW strength (in terms of the superwind parameter , as in S.-C. Yoon & M. Cantiello 2010), as well as subsequent shock-driven dynamical ejections (M. Clayton, 2018), the mass loss histories of the model sequences are coupled to an accelerated -law wind profile to construct spherically symmetric CSM density and column-density profiles that are directly compared to multi-epoch observations of SNe 1998S, 2005ip, 2017hcc, 2020ywx, and 2023ixf. We find decent agreement (within a factor of few) between the observationally inferred values and our model results for M⊙ and which give enhanced mass loss rates ( M⊙ yr-1) in the final decades to centuries before explosion. Such late phase heavy mass loss produces sufficiently dense CSM close to the SN progenitor to explain early-time observations (e.g. using flash-spectroscopy, X-rays and radio wavelengths) as well as late epoch estimates of X-ray column densities (Section 5). Our results quantitatively demonstrate that single star evolutionary pathways with pulsation-driven mass loss and consequent dynamical ejections can account for the observed diversity in the properties of Type II SN environments.
In future, missing physics such as effects of varying rotation and metallicity as well as binary interaction needs to be incorporated to improve the realism of our models, all of which are expected to have broader implications both for massive star evolution as well as their final fates in terms of the nature of the eventual SN. Also, the pulsation properties of other classes of massive stars such as WR stars (L. Grassitelli et al., 2016, 2018) and LBVs (J. A. Guzik et al., 1999; S. Justham et al., 2014) deserve investigation as they are supposed to be prime candidates for progenitors of both Type Ib/c and superluminous SNe. In addition, further improvements in our understanding of the interplay of radial pulsation and convective motions in RSG envelope through multidimensional simulations (J. A. Goldberg et al., 2022, 2025) and its role in driving mass loss are critical towards advances in the predictive power of evolutionary models by incorporation self-consistent treatment of stellar winds through physical mechanisms such as pulsations and shocks, as compared to the empirical approach adopted in this study and extant works in the literature (S.-C. Yoon & M. Cantiello, 2010; Moriya, Takashi J. & Langer, Norbert, 2015; V. A. Bronner et al., 2025). This will have important implications for SN modeling at large and their impact on cosmic evolution and dust formation ranging from our local environment to the high redshift universe in the era of large-scale surveys like the JWST and the LSST.
@@ 8 Acknowledgments
The authors acknowledge the benevolent support of the Department of Science and Technology (DST), Government of India. We thank Dr. Heloise Stevance for constructive suggestions and helpful comments on this work and Mr. Raphael Baer-Way for private communication and sharing data from R. Baer-Way et al. (2025). This research has made use of the High Performance Computing (HPC) resources 222https://www.iiap.res.in/centers/main-campus/computing-and-i-t/ made available by the Computer Center of the Indian Institute of Astrophysics (IIA), Bangalore.
@@ Appendix A Choice of parameter for accelerated wind
As described in Section 4.3.1, due to lack of a better understanding of the driving mechanism for the RSG wind out to larger distances few stellar radii where dust can form to drag the gas along, we choose to parameterize this kinematic process with the -velocity law as given by Equation (4). The normalized velocity profiles using this velocity law for various choices of commonly used in the literature is shown in Figure 11 which illustrates the change in the shape of the velocity profile, from convex to concave, very close to the stellar surface i.e. at , with increasing values of . Comparing the shape of the velocity curves with that of observations of -Veloris (P. D. Bennett, 2010), we select for the purpose of this study, as lower values of result in steeper velocity profiles inconsistent with the observed values, while higher values of do not match the shape of the observed velocity profiles close to the stellar surface.
@@ Appendix B Uncertainty estimates for calculating CSM densities
The uncertainties in the values of CSM densities inferred from X-ray mass loss rates and terminal wind velocities of SNe 2023ixf (W. V. Jacobson-Galán et al., 2025) and 2020ywx (R. Baer-Way et al., 2025), according to Equation (5), have been calculated by propagating the individual uncertainties in the estimation of these two properties from measurements of X-ray fluxes and fitting of spectral features, respectively. There are also additional uncertainties that involve assumptions regarding degree of ionization in the shock propagating through the CSM post explosion as well as the composition of the CSM but a quantitative estimate of these uncertainties are beyond the scope of this work.

### v3

@@ 6.4 Implications for pre-explosion dust formation around RSGs
RSG stars are often associated with a dusty circumstellar medium (walmswell_2012; clayton_2009; cannon_2021; niu_2023; gaston_2025). Yet the modest mass loss rates ( 10-6 M⊙) characterizing RSG winds are found unsuitable to synthesize dust grains from the gas-phase atoms (cherchneff2013b). Dust formation requires moderately low temperatures while the gas densities are still high to allow series of collisions leading to the formation of gas-phase precursors (sarangi2018book). In the RSG with a dense CSM, the enhanced mass loss can aid the formation of chemical compounds in the wind, depending on the wind velocities. As we have shown, each stellar model evolves in unique routes, which will result in a unique distribution of molecules and dust grains. If the enhanced mass loss is triggered during the He-burning phase or early C-burning phase, the dust formed in that gas will fly far from the star by the time the star goes to SN. This dust may already integrate with the ambient interstellar medium (ISM), and enhance the dust budget in the galaxy. Moreover, given the large distance, they are shielded from the strong ionizing SN flashwave at the point of core-collapse. However, this dust is difficult to detect in stellar spectroscopy in optical or IR, given it will be at a large distance (hence causing no significant extinction to the stellar light curve) and cold temperatures (making it indistinguishable with the ambient ISM).
On the other hand, in the RSGs with enhanced mass loss at the very end of their lives, the dusty CSM significantly impacts the stellar light curve and its post-explosion IR signatures. In addition, the composition of the CSM may also be altered based on the nature of the stellar evolution and mass loss. Even though the CSM is often identical in composition with the H-rich stellar envelope, a small variation in the abundance of the metals will impact the composition and dust masses. The formation of dust precursors in the gas is often controlled by some bottleneck gas-phase reactions (sar13). On the contrary, the growth of grains, which decides the grain size distributions, is proportional to time (hirashita_2012). The population of larger grains are more resilient to destruction processes post SN explosion, or shocks in the ISM (micelotta2018). In that regard, the timing of the enhanced mass loss plays key role in determining the grain sizes.
@@ 6.4.1 Numerical and computational challenges in obtaining pre-SN yields
In order to predict pre-explosion yields to obtain dust masses, our MESA runs need to be evolved through explosive oxygen and silicon burning that occurs on day-hour timescales. As already discussed, accurate energy generation requires the use of large nuclear networks (e.g. mesa_128.net) with implicit hydrodynamics and timesteps causing severe slowdowns (with a single run requiring more than a week on a -core CPU machine), making a similar systematic exploration of our parameter space in and , prohibitively expensive. A hybrid computing strategy – involving zone-by-zone parallelism on a graphics processing unit (GPU) by offloading the nuclear network calculations that form a major bottleneck of MESA with large nuclear networks – seems a promising approach to significantly reduce the large computational cost of evolving our runs through to CC. With an initial grid of such models, the recently proposed Nuclear Neural Network (NNN) framework of Grichener_2025 can be employed to further refine the grid that seems essential to study stochastic effects of convection, shell merging etc. on final yields and associated dust masses.
@@ 7 Conclusions
This work presents a comprehensive, end-to-end investigation of how pulsating RSGs can create a dense, possibly dusty, CSM through enhanced mass loss episodes in the final years before CC, which shapes the multi-wavelength appearance of the ensuing SN explosion. From a dense grid of MESA models of varying M⊙ and PDSW strength (in terms of the superwind parameter , as in 2010ApJ...717L..62Y), as well as subsequent shock-driven dynamical ejections (clayton2018a), the mass loss histories of the model sequences are coupled to an accelerated -law wind profile to construct spherically symmetric CSM density and column-density profiles that are directly compared to multi-epoch observations of SNe 1998S, 2005ip, 2017hcc, 2020ywx, and 2023ixf. We find decent agreement (within a factor of few) between the observationally inferred values and our model results for M⊙ and which give enhanced mass loss rates ( M⊙ yr-1) in the final decades to centuries before explosion. Such late phase heavy mass loss produces sufficiently dense CSM close to the SN progenitor to explain early-time observations (e.g. using flash-spectroscopy, X-rays and radio wavelengths) as well as late epoch estimates of X-ray column densities (Section 5). Our results quantitatively demonstrate that single star evolutionary pathways with pulsation-driven mass loss and consequent dynamical ejections can account for the observed diversity in the properties of Type II SN environments.
In future, missing physics such as effects of varying rotation and metallicity as well as binary interaction needs to be incorporated to improve the realism of our models, all of which are expected to have broader implications both for massive star evolution as well as their final fates in terms of the nature of the eventual SN. Also, the pulsation properties of other classes of massive stars such as WR stars (2016A&A...590A..12G; 2018A&A...614A..86G) and LBVs (1999LNP...523..337G; 2014ApJ...796..121J) deserve investigation as they are supposed to be prime candidates for progenitors of both Type Ib/c and superluminous SNe. In addition, further improvements in our understanding of the interplay of radial pulsation and convective motions in RSG envelope through multidimensional simulations (2022ApJ...929..156G; 2025A&A...703A..61B; 2025arXiv251014875M) and its role in driving mass loss are critical towards advances in the predictive power of evolutionary models by incorporation self-consistent treatment of stellar winds through physical mechanisms such as pulsations and shocks, as compared to the empirical approach adopted in this study and extant works in the literature (2010ApJ...717L..62Y; ml2015; 2025A&A...703A..61B; 2025MNRAS.543.3929S). This will have important implications for SN modeling at large and their impact on cosmic evolution and dust formation ranging from our local environment to the high redshift universe in the era of large-scale surveys like the JWST and the LSST.
@@ 8 Acknowledgments
The authors acknowledge the support of the Department of Science and Technology (DST), Government of India. We thank Dr. Heloise Stevance and Professor Jan Eldridge for constructive suggestions and helpful comments on this work, and Mr. Raphael Baer-Way for private communication and for sharing data from Baer-Way_2025. This research has made use of the High Performance Computing (HPC) resources222https://www.iiap.res.in/centers/main-campus/computing-and-i-t/ made available by the Computer Center of the Indian Institute of Astrophysics (IIA), Bangalore.
@@ Appendix A Choice of parameter for accelerated wind
As described in Section 4.3.1, due to lack of a better understanding of the driving mechanism for the RSG wind out to larger distances few stellar radii where dust can form to drag the gas along, we choose to parameterize this kinematic process with the -velocity law as given by Equation (4). The normalized velocity profiles using this velocity law for various choices of commonly used in the literature is shown in Figure 11 which illustrates the change in the shape of the velocity profile, from convex to concave, very close to the stellar surface i.e. at , with increasing values of . Comparing the shape of the velocity curves with that of observations of -Veloris (2010ASPC..425..181B), we select for the purpose of this study, as lower values of result in steeper velocity profiles inconsistent with the observed values, while higher values of do not match the shape of the observed velocity profiles close to the stellar surface.
@@ Appendix B Uncertainty estimates for calculating CSM densities
The uncertainties in the values of CSM densities inferred from X-ray mass loss rates and terminal wind velocities of SNe 2023ixf (2025arXiv250811747J) and 2020ywx (Baer-Way_2025), according to Equation (5), have been calculated by propagating the individual uncertainties in the estimation of these two properties from measurements of X-ray fluxes and fitting of spectral features, respectively. There are also additional uncertainties that involve assumptions regarding degree of ionization in the shock propagating through the CSM post explosion as well as the composition of the CSM but a quantitative estimate of these uncertainties are beyond the scope of this work.